• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name brads

    Log entry time 12:12:44 on May 27, 2008

    Entry number 239880

    keyword=Controlled Access activity

    Spent some time troubleshooting the BB Cerenkov and chasing down some bad signals. There seem to be some bad patch cables running from the PMTs to the front-end. Swapped to differenct cable for PMT 14, have to bring another bundle of 60ns cables from Testlab for next access.

    A couple of the summing modules for PMTs 11-20 appear to have a lot of 60Hz noise on them. Have to swap them out next access...

    Adjusted the cerenkov group discriminator threshold to 100mV to get it above some of the 60Hz ripple from the bad summing modules. Now that I think about it, that discriminator is remotely controllable so I just bumped the offset up by +50mV. Damn. I updated the offset in '/adaqfs/home/adaq/brad/camac_control/threshold' so the software reports the correct threshold...

    Bumped the camera pointing to the BB Cerenkov gas panel so I had to ramp BB down, reset the camera, and ramp BB back up. (BB needs to be ramped down because the handheld TV used to view the camera while moving/focusing it gets creamed by the the BB field.)

    Bob Michaels also did some work in the RHRS hut (a crate needed for his parasitic test died).